# Robust Pronoun Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

### v2

@@ 10 Related Work
Pronoun fidelity.Hossain et al. (2023) and Ovalle et al. (2023a) both study pronoun fidelity when models are prompted with a pronoun series to use for an individual, but they only consider simplistic pronoun use with no more than one person at a time. Although we look at within-language pronoun use, faithful pronoun use in context has also been studied in machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023), where there is also a ground truth. Similar to our work, Sharma et al. (2022) injects context with an explicit coreference to encourage faithful pronoun translation. However, none of these papers explore the robustness of pronoun fidelity in the presence of distractors.
Reasoning with pronouns. Most existing work about LLM reasoning with pronouns focuses on the task of coreference resolution, i.e., the ability to identify the connection between a pronoun and an entity, which may not translate to faithful reuse of that pronoun later, as in our work. Reasoning with pronouns typically uses Winograd schemas (Levesque et al., 2012; Abdou et al., 2020; Emelin and Sennrich, 2021), or Winograd-like schemas about named individuals (Webster et al., 2018; Zhao et al., 2018), or people referred to by their occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include singular they (Baumler and Rudinger, 2022) and with neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023), as we do.
@@ 11 Conclusion
We introduce the task of pronoun fidelity to evaluate robust, faithful and harm-free pronoun use in language models, and we present RUFF, a dataset we designed to evaluate it. We find evidence of faithful pronoun use only in a very simple setting, i.e., where only one person is discussed. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Even adding a single sentence about a second individual with a different pronoun causes accuracy to drop dramatically, showing that pronoun fidelity is neither robust to non-adversarial distractors nor due to “reasoning.” As more distractor sentences are added, encoder-only models perform better overall, but increasingly revert to biased predictions, while decoder-only models get increasingly distracted. Our results show that in a setting that is very simple for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers to bridge the performance gaps we report and to more carefully evaluate “reasoning,” especially when simple repetition could inflate perceptions of model performance.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, as well as Aaron Mueller, Marius Mosbach, Vlad Niculae and Yanai Elazar, for feedback and discussions about math, plots and framing. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Limitations
Shallow heuristics. Much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019; Elazar et al., 2021). We caution readers that our dataset also gives a very generous estimate of model reasoning performance for multiple reasons. Many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. The association between janitor and mop is strong compared to child (the participant paired with janitor) and mop, and could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.” Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset in its current form should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
Whose bias? Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc.
Appendix B List of Occupations
@@
Appendix C Context Template Construction
Semantically bleached templates such as He is an accountant are well-established in the literature for testing word embedding associations (Caliskan et al., 2017), but it is unnatural to use more than one consecutively, which is necessary to evaluate robustness in this setting. On the other hand, natural corpora like Levy et al. (2021) contain occupation-specific sentences that cannot be used more generally for a controlled setting across occupations and participants. As a compromise between a controlled setting and naturalistic data, we create context templates with generic themes that apply to all humans, e.g., universal emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.).
For each case, we create 10 explicit templates which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. The first two sentences of the context (i.e., the introduction and the first distractor) are always sampled from the explicit templates, and the rest are sampled from the implicit templates, as this reflects natural and coherent use of pronouns in discourse.
In both cases, we create five templates with terms with positive connotations (e.g., full, happy) and five opposing templates (i.e., hungry, unhappy). We denote $exp\_pos_{i}$ as the i-th positive explicit template where i ranges from 1 to 5, and $exp\_neg_{i}$ is the corresponding negative version. The introduction template can be selected from any of these 10 possibilities and filled with one of four pronouns.
@@ Appendix C Context Template Construction
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{5}\}$. For subsequent distractor templates, we sample without replacement from these implicit templates.
Appendix D Annotator Demographics
@@
Appendix E Annotation Instructions
E.1 Task 1 Description
@@ E.1.1 Grammaticality
E.1.2 Questions about pronouns
@@ E.2 Task 2 Description
Appendix F Experimental Details
@@ G.2 Chain-of-thought prompting
Appendix H Context-free pronoun predictions by model
@@
Appendix I Results with Vanilla and Chain-of-Thought Prompting
I.1 Vanilla prompting
Prompting is a different model evaluation mechanism than log likelihoods, with higher task demands that lead to lower performance than log likelihoods with both base models and instruction fine-tuned chat models (Hu and Levy, 2023; Hu and Frank, 2024; Kauf et al., 2024). We thus expect vanilla prompting results (using the prompts listed in Appendix G) to be worse than results with log likelihoods. Indeed, Figure 9(a) shows that Llama-2-chat prompting performance is lower than Llama-2 evaluated with log likelihoods, even with no distractors. Figure 9(b) shows the results of standard prompting with FLAN-T5, an encoder-decoder model which shows similar patterns of degradation to decoder-only models. Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots.
@@ I.1 Vanilla prompting
I.2 Chain-of-thought prompting

### v3

@@ 10 Related Work
Pronoun fidelity.Hossain et al. (2023) and Ovalle et al. (2023) both study pronoun fidelity when models are prompted with a pronoun series to use for an individual, but they only consider simplistic pronoun use with no more than one person at a time. Although we look at within-language pronoun use, faithful pronoun use in context has also been studied in machine translation (Müller et al., 2018; Voita et al., 2018; Fernandes et al., 2023), where there is also a ground truth. Similar to our work, Sharma et al. (2022) injects context with an explicit coreference to encourage faithful pronoun translation. However, none of these papers explore the robustness of pronoun fidelity in the presence of distractors.
Reasoning with pronouns. Most existing work about LLM reasoning with pronouns focuses on the task of coreference resolution, i.e., the ability to identify the connection between a pronoun and an entity, which may not translate to faithful reuse of that pronoun later, as in our work. Reasoning with pronouns typically uses Winograd schemas (Levesque et al., 2012; Abdou et al., 2020; Emelin and Sennrich, 2021), or Winograd-like schemas about named individuals (Webster et al., 2018; Zhao et al., 2018), or people referred to by their occupation (Rudinger et al., 2018; Levy et al., 2021). Most studies focus on he and she, but recent work has expanded to include singular they (Baumler and Rudinger, 2022) and neopronouns (Cao and Daumé III, 2021; Felkner et al., 2023), as we do.
@@ 11 Conclusion
We introduce the task of pronoun fidelity to evaluate robust, faithful and harm-free pronoun use in language models, and we present RUFF, a dataset we designed to evaluate it. We find evidence of faithful pronoun use only in a very simple setting, i.e., when only one person is discussed. Even here, models show significant performance disparities with neopronouns, singular they and she/her/her, compared to he/him/his. Even adding a single sentence about a second individual with a different pronoun causes accuracy to drop dramatically, showing that pronoun fidelity is neither robust to non-adversarial distractors nor due to “reasoning.” As more distractor sentences are added, encoder-only models perform better overall, but increasingly revert to biased predictions, while decoder-only models get increasingly distracted. Our results show that in a setting that is very simple for humans, widely-used large language models are unable to robustly and faithfully reason about pronouns, and continue to amplify discrimination against users of certain pronouns. We encourage researchers to bridge the performance gaps we report and to more carefully evaluate “reasoning,” especially when simple repetition could inflate perceptions of model performance.
12 Limitations
Shallow heuristics. Much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019; Elazar et al., 2021). We caution readers that our dataset also gives a very generous estimate of model reasoning performance, as many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. Janitor is more strongly associated with mop than child, which could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.” Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset in its current form should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
Whose bias? Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc. In addition, pronoun fidelity is only one dimension of inclusive language model behaviour, and indeed only one way in which misgendering occurs in language, even in morphologically poor languages such as English.
@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, as well as Aaron Mueller, Marius Mosbach, Vlad Niculae, Yanai Elazar, our action editor Hai Zhao, and our anonymous TACL reviewers, for feedback on math, plots and framing, that improved this work. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. Anne Lauscher’s work is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A List of Occupations
@@
Appendix B Context Template Construction
For each grammatical case, we create 10 explicit templates, which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. An introduction and the first distractor are always sampled from the explicit templates, and the rest are sampled from the implicit templates, as this reflects natural and coherent use of pronouns in discourse.
For both the explicit and implicit cases, we create five templates with terms with positive connotations (e.g., full, happy) and five templates with the opposite polarity (i.e., hungry, unhappy). We denote $exp\_pos_{i}$ as the i-th positive explicit template where i ranges from 1 to 5; $exp\_neg_{i}$ is the corresponding negative version. The introduction template can be selected from any of these 10 possibilities and filled with one of four pronouns.
@@ Appendix B Context Template Construction
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{3},imp\_neg_{5}\}$. For subsequent distractor templates, we sample without replacement from these implicit templates.
Appendix C Annotator Demographics
@@
Appendix D Annotation Instructions
D.1 Task 1 Description
@@ D.1.1 Grammaticality
D.1.2 Questions about pronouns
@@ D.2 Task 2 Description
Appendix E Experimental Details
@@ Appendix F Prompting
F.1 Vanilla prompting
F.2 Chain-of-thought prompting
Appendix G Context-free pronoun predictions by model
@@
Appendix H Results with Vanilla and Chain-of-Thought Prompting
H.1 Vanilla prompting
Prompting is a different model evaluation mechanism than log likelihoods, with higher task demands that lead to lower performance than log likelihoods with both base models and instruction fine-tuned chat models (Hu and Levy, 2023; Hu and Frank, 2024; Kauf et al., 2024). We thus expect vanilla prompting results (using the prompts listed in Appendix F) to be worse than results with log likelihoods. Indeed, Figure 11(a) shows that Llama-2-chat prompting performance is lower than Llama-2 evaluated with log likelihoods, even with no distractors. Figure 11(b) shows the results of standard prompting with FLAN-T5, an encoder-decoder model which shows similar patterns of degradation to decoder-only models. Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots.
@@ H.1 Vanilla prompting
H.2 Chain-of-thought prompting